a bill of 13.75 was paid in half dollars and quarters. the number of half dollars exceeded the number of quarters by 5. how many of each are there
Answer by solver91311(12126)

number of half dollars: h
number of quarters: q

value of h half dollars: 50h cents
value of q quarters: 25q cents

Total money: $13.75, which can be expressed as 1375 cents, so:

50h + 25q = 1375

5 more half dollars than quarters: h = q + 5, so:

50(q + 5) + 25q = 1375

Can you solve it from here or do you need more help?